# All-in-One Video Restoration under Smoothly Evolving Unknown Weather Degradations

### v1

@@ Abstract
All-in-one image restoration aims to recover clean images from diverse unknown degradations using a single model. But extending this task to videos faces unique challenges. Existing approaches primarily focus on frame-wise degradation variation, overlooking the temporal continuity that naturally exists in real-world degradation processes. In practice, degradation types and intensities evolve smoothly over time, and multiple degradations may coexist or transition gradually. In this paper, we introduce the Smoothly Evolving Unknown Degradations (SEUD) scenario, where both the active degradation set and degradation intensity change continuously over time. To support this scenario, we design a flexible synthesis pipeline that generates temporally coherent videos with single, compound, and evolving degradations. To address the challenges in the SEUD scenario, we propose an all-in-One Recurrent Conditional and Adaptive prompting Network (ORCANet). First, a Coarse Intensity Estimation Dehazing (CIED) module estimates haze intensity using physical priors and provides coarse dehazed features as initialization. Second, a Flow Prompt Generation (FPG) module extracts degradation features. FPG generates both static prompts that capture segment-level degradation types and dynamic prompts that adapt to frame-level intensity variations. Furthermore, a label-aware supervision mechanism improves the discriminability of static prompt representations under different degradations. Extensive experiments show that ORCANet achieves superior restoration quality, temporal consistency, and robustness over image and video-based baselines. Code is available at https://github.com/Friskknight/ORCANet-SEUD.
@@ I Introduction
However, there are three limitations in previous work on all-in-one restoration for videos. (1) Limited modeling of temporal correlation of degradation in TUD. In practice, factors such as rain density, haze thickness, and snow motion change smoothly over time, and different weather types may coexist or transition gradually. In contrast, the TUD setting [31] applies degradations to each frame independently without temporal linkage. This design neglects the physical continuity of real-world degradations. This design breaks the natural continuity of degradation and produces unrealistic temporal changes. (2) Limited modeling of cross-frame dependencies in prompt design. Prompt-based AiOIR methods mainly focus on single-image restoration. Textual and multimodal prompts require precise external inputs and face cross-modal alignment issues and model complexity [32, 33, 34, 35, 36]. Visual prompts extract embeddings from the degraded image, but they are usually generated per frame and remain independent across time [4, 37, 38]. As a result, they fail to capture how degradations evolve. (3) Limited modeling of compound degradations. Many all-in-one restoration methods struggle to handle co-occurring degradations. For example, PromptIR [37] and PIP [4] do not include mechanisms to prevent similarity or conflict among task prompts. T3DiffWeather [36] uses a predefined prompt pool, which limits coverage of unseen or mixed degradations. TAP [39] models task relations with low-rank and contrastive constraints but relies on predefined degradation correlations, which limits flexibility in complex or dynamic scenarios.
@@ IV-A Dataset and Evaluation Metrics
Figure 5: Recovery results of single-type degraded videos in setting 1.
@@ IV-B Evaluation on Dynamic and Compound Degradation
Figure 6: Recovery results of single-type degraded videos in setting 2.
@@
Figure 7: Recovery results of single-type degraded videos in setting 4.
@@ IV-C Degradation Frame Stability and Consistency
IV-D Ablation Studies
@@ V Conclusion
We also propose ORCANet, a unified restoration framework that integrates coarse intensity–aware dehazing, flow-guided bidirectional propagation, and a two-level prompt design that captures both segment-level degradation identity and frame-level variation. Extensive experiments show that ORCANet achieves superior performance across a wide range of dynamic and compound degradation conditions. The model produces stable and temporally consistent results and demonstrates strong generalization when degradation strength and type change over time. These results confirm the importance of modeling temporal continuity and degradation-aware prompts in video all-in-one restoration. Future work will further broaden the scope of degradation types, and develop lightweight variants for real-time applications.

### v2

@@ Abstract
All-in-one image restoration aims to recover clean images from diverse unknown degradations using a single model. But extending this task to videos faces unique challenges. Existing approaches primarily focus on frame-wise degradation variation, overlooking the temporal continuity that naturally exists in real-world degradation processes. In real world, degradation types and intensities evolve smoothly over time, and multiple degradations may coexist or transition gradually. In this paper, we introduce the Smoothly Evolving Unknown Degradations (SEUD) scenario to simulate such real-world processes, where both the active degradation set and degradation intensity change continuously over time. To support this scenario, we design a flexible synthesis pipeline that generates temporally coherent videos with single, compound, and evolving degradations. To address the challenges in the SEUD scenario, we propose an all-in-One Recurrent Conditional and Adaptive prompting Network (ORCANet). First, a Coarse Intensity Estimation Dehazing (CIED) module estimates haze intensity using physical priors and provides coarse dehazed features as initialization. Second, a Flow Prompt Generation (FPG) module extracts degradation features. FPG generates both static prompts that capture segment-level degradation types and dynamic prompts that adapt to frame-level intensity variations. Furthermore, a label-aware supervision mechanism improves the discriminability of static prompt representations under different degradations. Extensive experiments show that ORCANet achieves superior restoration quality, temporal consistency, and robustness over image and video-based baselines. Code is available at https://github.com/Friskknight/ORCANet-SEUD.
@@ I Introduction
However, there are three limitations in previous work on all-in-one restoration for videos. (1) Limited modeling of temporal correlation of degradation in TUD. In practice, factors such as rain density, haze thickness, and snow motion change smoothly over time, and different weather types may coexist or transition gradually. In contrast, the TUD setting [31] applies degradations to each frame independently without temporal linkage. This design neglects the physical continuity of real-world degradations. This design breaks the natural continuity of degradation and produces unrealistic temporal changes. (2) Limited modeling of cross-frame dependencies in prompt design. Prompt-based AiOIR methods mainly focus on single-image restoration. Textual and multimodal prompts require precise external inputs and face cross-modal alignment issues and model complexity [32, 33, 34, 35, 36]. Visual prompts extract embeddings from the degraded image; but they are usually generated per frame and remain independent across time [4, 37, 38]. As a result, they fail to capture how degradations evolve. (3) Limited modeling of compound degradations. Many all-in-one restoration methods struggle to handle co-occurring degradations. For example, PromptIR [37] and PIP [4] do not include mechanisms to prevent similarity or conflict among task prompts. T3DiffWeather [36] uses a predefined prompt pool, which limits coverage of unseen or mixed degradations. TAP [39] models task relations with low-rank and contrastive constraints but relies on predefined degradation correlations, which limits flexibility in complex or dynamic scenarios.
@@ IV-A Dataset and Evaluation Metrics
Figure 5: Comparison of visual restoration results of rainy degraded video frame.
@@ IV-B Evaluation on Dynamic and Compound Degradation
Figure 6: Comparison of visual restoration results of foggy degraded video frame.
@@
Figure 7: Comparison of visual restoration results under compound degraded video frame.
@@ IV-C Degradation Frame Stability and Consistency
IV-D Real-World Video Restoration Results
Figure 11: Visual comparisons on real-world video restoration results from GoProHazy [60] and NTURain [61]. Columns (a)–(e) show the input and the results of PromptIR, BasicVSR++, T3DiffWeather, and ORCANet. Please zoom in for clearer view.
We evaluate real-world restoration on videos from GoProHazy [60] and NTURain [61]. We compare the visual results with PromptIR, BasicVSR++, and T3DiffWeather. For PromptIR and T3DiffWeather, we test two settings. One setting uses the official pretrained weights. The other setting uses weights that we train on SEUD. We report the better result for each method. Fig. 11 shows the visual comparisons. On GoProHazy, BasicVSR++ and T3DiffWeather leave visible haze in some local regions. PromptIR removes haze well, but it introduces noticeable noise in several areas. ORCANet restores the nearby regions more uniformly and it preserves local details across the frame. On NTURain, the videos contain many thin rain streaks, and some scenes also include haze. T3DiffWeather recovers richer textures, but it keeps most rain traces. BasicVSR++ removes many rain streaks, but it leaves residual streaks around complex regions, and haze removal is still uneven. ORCANet handles these compound degradations better and it produces the cleanest and most stable appearance in our comparisons.
IV-E Ablation Studies
@@ V Conclusion
We also propose ORCANet, an all-in-one restoration framework that integrates coarse intensity–aware dehazing, flow-guided bidirectional propagation, and a two-level prompt design that captures both segment-level degradation identity and frame-level variation. Extensive experiments show that ORCANet achieves superior performance across a wide range of dynamic and compound degradation conditions. The model produces stable and temporally consistent results and demonstrates strong generalization when degradation strength and type change over time. These results confirm the importance of modeling temporal continuity and degradation-aware prompts in video all-in-one restoration. Future work will further broaden the scope of degradation types, and develop lightweight variants for real-time applications.